# Strategic Investment to Mitigate Transition Risks

### v1

@@ Abstract
This paper investigates strategic investments needed to mitigate transition risks, particularly focusing on sectors significantly impacted by the shift to a low-carbon economy. It emphasizes the importance of tailored sector-specific strategies and the role of government interventions, such as carbon taxes and subsidies, in shaping corporate behavior. In providing a multi-period framework, this paper evaluates the economic and operational trade-offs companies face under four various decarbonization scenarios—immediate, quick, slow, and no transitions. The analysis provides practical insights for both policymakers and business leaders, demonstrating how regulatory frameworks and strategic investments can be aligned to manage transition risks while optimizing long-term sustainability effectively. The findings contribute to a deeper understanding of the economic impacts of regulatory policies and offer a comprehensive framework to navigate the complexities of transitioning to a low-carbon economy.
@@ 1 Introduction
1.1 Climate Risks
Mitigating climate change requires rapid and extensive decarbonization of the global economy. Climate change poses severe risks, such as altering extreme weather patterns and threatening ecosystems. Nonlinear climate dynamics can lead to sudden, irreversible changes, such as a polar ice melt or an ecosystem collapse, representing critical tipping points in Earth’s systems (Lenton et al. (2008)). The 2015 Paris Agreement underscores the urgency of limiting a global temperature rise to well below 2°C to prevent catastrophic outcomes, with net-zero carbon emissions needed by mid-century to meet this goal (Holden et al. (2018)).
@@ 1.1 Climate Risks
Achieving decarbonization demands transformative changes across industries. While sectors like a power generation have made strides toward a renewable energy adoption, industries such as a steel and aviation lag due to underdeveloped low-carbon alternatives (Kober et al. (2020), Dietz et al. (2020)). These challenges highlight the importance of understanding climate risks, which fall into two categories: physical and transition risks. Physical risks encompass direct impacts of climate change, such as extreme weather events, while transition risks arise from shifting to a low-carbon economy driven by regulations, technologies, and societal demands (Carney (2015)). This paper prioritizes transition risks, as they offer more scope for a strategic intervention and remain underexplored in quantitative modeling (Cisagara (2024)).
1.2 Understanding Transition Risk
@@
Transition risks result from structural changes in industries due to the shift toward a low-carbon economy. High-emission sectors, such as steel, aviation, and energy-intensive manufacturing, face challenges adapting to stricter regulations and cleaner technologies. Policy uncertainty adds complexity, creating financial risks for firms and investors (Holden et al. (2018)). These risks include stranded assets, increased credit risk, and volatility in asset returns (Semieniuk et al. (2021)).
@@
While low-carbon policies, such as carbon taxes and emissions trading schemes, are essential for meeting climate objectives, their implementation introduces non-trivial transitional challenges (Bank of England (2018)). Companies are now evaluated not only on their financial metrics but also on their resilience to climate risks (Schulte and Hallstedt (2018)). Addressing these risks effectively requires targeted strategies and robust quantitative models.
1.3 Low-Carbon Transformation: Balancing Benefits and Challenges
@@
The shift toward a low-carbon economy offers opportunities, such as cost savings, market leadership, and innovation, alongside significant challenges. Companies like Tesla and BP have successfully aligned with low-carbon strategies, reaping benefits such as new revenue streams and enhanced competitiveness. However, this effort of transition is capital-intensive, with high upfront costs and regulatory uncertainty posing risks to companies (NGFS (2019), Schulte and Hallstedt (2018)). Additionally, the rapid pace of technological advancements can render current investments obsolete, further complicating long-term planning.
@@
Policymakers and businesses must balance a climate action with an economic stability. In the next section, we introduce a stylized model to guide low-carbon transition strategies, helping industries navigate uncertainties and optimize their transformation efforts.
2 Preliminaries and Model Building
@@ 2.1 Motivations
2.1.1 Sector-Specific Strategies and Investments in the Low-Carbon Transition
Different industries face unique challenges and investment needs in the low-carbon transition. Key sectors include:
  • Energy: Transitioning to renewable energy requires upgrading infrastructure and advancing storage technologies (Boyd (2013)).
  • Transportation: Decarbonization efforts focus on electric vehicles, hydrogen fuel, and charging networks (Tyfield (2013)).
  • Manufacturing: Industries such as steel and chemical processing are exploring low-carbon processes like hydrogen-based production and carbon capture technologies.
  • Agriculture: Investments in regenerative practices and ecosystem preservation are critical for reducing emissions (van Veelen (2021)).
Transition costs are substantial. Estimates suggest global spending on net-zero initiatives could reach \$275 trillion by 2050, with varying investment levels across sectors (McKinsey (2021)). Tailored, sector-specific strategies are crucial to balancing costs and achieving decarbonization goals effectively.
2.1.2 Balancing Costs and Benefits of Low-Carbon Transition
Investing in the low-carbon transition confers great benefits across various industries by driving long-term sustainability, reducing regulatory risks, and enhancing competitive positioning. Reducing carbon emissions can lead to cost savings through improved energy efficiency and reduced fuel expenses for energy-intensive industries such as manufacturing and transportation. In the financial sector, supporting low-carbon initiatives can mitigate portfolio risks linked to climate change, while at the same time also opening new investment opportunities in green technologies and sustainable businesses. Industries such as real estate and construction benefit by developing energy-efficient buildings that attract environmentally conscious tenants, boosting property value. Additionally, adopting low-carbon strategies can enhance the brand reputation for corporate responsibility and environmental stewardship, meet evolving consumer preferences for sustainability, and ensure compliance with future climate regulations, positioning companies for growth in a low-carbon economy. However, the upfront costs of adopting cleaner technologies can strain businesses, particularly if the returns on investment are not immediate. For example, investing in renewable energy infrastructure may tie up capital that could otherwise be used for business expansion or other needs.
@@ 2.1.2 Balancing Costs and Benefits of Low-Carbon Transition
Given these considerations, this paper proposes a stylized analytical framework to determine the optimal low-carbon transition strategy across different sectors. The framework aims at capturing the unique underlying dynamics and investment needs of each sector, helping industries navigate the complexities of transitioning toward a sustainable, low-carbon future.
2.2 Variables Descriptions
@@
The parameters and variables in this framework are introduced with the aim of replicating the dynamics of low-carbon transition investments and their effects on various industry operations, offering insights into investment strategies. Below, we provide a detailed description of each variable and parameter.
@@
Selling price per unit ($p$): This parameter represents the price at which the product of interest is sold in the market. Taking the selling price into consideration is to describe and quantify the revenue. In the context of low-carbon transition, businesses may find opportunities to increase prices for eco-friendly products due to an increased consumer demand for sustainability. For instance, Tesla’s Model 3 and other EVs are often priced higher than traditional internal combustion engine vehicles (ICEs), yet some consumers are willing to pay the premiums for the environmental advantages.
@@
Production Cost per unit ($c$): Production cost per unit includes all expenses incurred in manufacturing each unit, such as raw materials, labor, and energy. Analyzing production costs is essential to evaluating profitability. A study by Siddique and Sultana (2018) found that switching to sustainable sourcing practices initially raised production costs due to the adoption of cleaner and more expensive technologies, but long-term benefits such as energy savings and brand value will outweigh the initial investment.
@@
Selling Profit per unit ($p-c$): This parameter reflects the profit earned from each unit sold, determined by subtracting production costs from the selling price. It captures the direct profitability of the product after accounting for costs. Profit is the primary goal for a company when making decisions, as it ensures financial viability and supports long-term growth, influencing how much of the company’s revenue can be allocated toward investments, including those for sustainable transitions.
@@
Total Asset ($A$): Total asset represents all financial resources available to a company, encompassing both current and long-term assets. This metric is essential for evaluating the company’s capacity to invest in production and growth opportunities, providing insight into the scale and potential of the business. Research indicates that firms with larger asset bases are more inclined to invest significantly in sustainability initiatives, as they can support longer investment horizons and absorb higher upfront costs associated with such projects (Eccles et al. (2014)). For example, Amazon has invested \$2 billion in renewable energy through its Climate Pledge Fund, leveraging its extensive financial resources to lead in corporate sustainability (Amazon (2020)). Consequently, total assets not only reflect a company’s financial strength but also its commitment to sustainable practices and long-term value creation.
@@
Transition Investment Ratio ($\alpha$): The Transition Investment Ratio is a critical metric in our model, measuring the proportion of total assets allocated to low-carbon initiatives. This ratio is essential as it reflects a company’s commitment to sustainability and its responsiveness to market dynamics and regulatory pressures. By strategically adjusting this ratio, businesses can enhance their reputation and mitigate risks associated with climate-related regulations and consumer preferences. Our primary goal with this framework is to identify the optimal transition investment ratio that balances financial performance with sustainability objectives, ensuring that companies thrive economically while still contributing positively to the environment.
@@
Practically, the Transition Investment Ratio offers insights into how effectively a company aligns its financial resources with sustainability goals. A higher ratio signifies a greater investment in low-carbon initiatives, leading to long-term cost savings and competitive advantages. For instance, IKEA’s allocation of over 3 billion pounds toward renewable energy infrastructure exemplifies the benefits of a strategic adjustment in this ratio (IKEA (2020)). By investing in solar and wind energy projects, IKEA not only reduced its operational reliance on fossil fuels but also achieved improved operational efficiency and lower carbon emissions. Improved operational efficiency can manifest in reduced energy costs, as renewable energy sources often lead to lower utility bills. IKEA’s renewable energy infrastructure allows it to produce a substantial portion of its energy needs in a sustainable manner, ultimately lowering the company’s operational costs and enhancing its energy independence. Additionally, the transition to renewable energy directly impacts carbon emissions, as utilizing solar and wind energy decreases IKEA’s greenhouse gas footprint. This shift aligns with regulatory expectations and resonates with consumers who favor environmentally responsible brands. Thus, the Transition Investment Ratio is pivotal in our model, encapsulating the strategic allocation of resources toward sustainability and driving both operational efficiency and lower carbon emissions.
@@
As previously mentioned, undergoing a low-carbon transition and investing in this transition can decrease production costs and increase selling prices due to improved operational efficiency, stronger brand value, and an alignment with consumer preferences. However, in the current framework, we assume both the production costs and the selling prices to be constant for simplicity. This assumption allows us to isolate the decision-making process regarding the resource allocation and emphasize the role of the Transition Investment Ratio in balancing financial performance and sustainability objectives without introducing additional complexities. Future work could build on this framework by incorporating functional relationships between these parameters—such as production costs and selling prices—and the Transition Investment Ratio.
@@
Enterprise Low-carbon Production Efficiency Coefficient ($k$): This coefficient measures the improvement in the firm’s production efficiency resulting from low-carbon initiatives, with a value greater than 1 indicating a gain in efficiency due to cleaner technologies. In our framework, we assume a higher value of $k$ signifies greater production efficiency of the firm after investing in the transition to sustainable practices. For instance, Siemens reported a remarkable 20% increase in production efficiency after transitioning from traditional fossil fuel-based energy to renewable energy sources and incorporating energy-efficient manufacturing technologies (AG (2019)). This transition not only reduced their carbon footprint but also streamlined their operations, demonstrating that investments in sustainability can yield great returns in efficiency.
@@
Original Productivity Coefficient ($\beta$): The value of this coefficient ranges between 0 and 1. A lower value indicates that, given a fixed amount of raw materials, the firm’s production efficiency is relatively low, resulting in fewer products being produced per unit of input, which reflects suboptimal resource utilization. Conversely, a value closer to 1 suggests higher efficiency in material utilization, allowing the enterprise to produce more output with the same raw material input, indicating both optimized resource allocation and production process efficiency. Changes in this coefficient can thus reflect the current production efficiency of the enterprise and offer insights into potential improvements under low-carbon transition strategies. There is evidence to suggest that many firms that start with below-average efficiency have successfully turned their operations around after investing in cleaner technologies. Research in Weng et al. (2015) indicates that companies in the regions with lower initial productivity often experience the most significant gains when transitioning to more efficient, low-carbon systems. By enhancing productivity through sustainable practices, these firms not only improve their operational performance but also position themselves for greater profitability in the long run. Thus, the firms with lower $\beta$ values should prioritize upgrades to enhance productivity and better align with sustainability goals.
@@
Carbon Price ($B$): The carbon price signifies the cost of emissions imposed by regulatory frameworks and is essential to a company’s profitability. This price is sensitive to regulatory or policy changes, with high carbon prices creating significant financial pressure on firms to reduce emissions. When the market carbon price is elevated, it can substantially diminish a company’s profit margins, compelling businesses to reassess their operational strategies. In response to these challenges, companies are incentivized to invest in cleaner technologies and adopt sustainable practices to mitigate their carbon liabilities. As firms strive to maintain profitability while adhering to environmental regulations, understanding and managing the implications of carbon pricing becomes crucial for their long-term success.
@@
Carbon Intensity ($I$): Carbon intensity refers to the amount of carbon dioxide (CO2) emissions produced per unit of output, typically measured in terms of emissions per product manufactured or per unit of energy consumed. It serves as a critical metric for companies transitioning to low-carbon production, as reducing carbon intensity is essential for minimizing environmental impact while maintaining operational efficiency.
@@
In the literature, carbon intensity is usually defined as a measure of the environmental performance of a company’s production processes. For instance, the IPCC emphasizes that carbon intensity is a key indicator of how effectively an organization utilizes resources and minimizes emissions in its operations (Change et al. (2014)). Similarly, the World Resources Institute (WRI) defines carbon intensity as a means to assess the relationship between output and emissions, allowing for comparisons across industries and over time (Institute (2015)). In addition, carbon intensity is defined more specifically as the ratio of expected annual emissions to expected revenue in Fang et al. (2019). This formulation highlights the firm’s contribution to climate change per unit of revenue, thus linking environmental impact directly to economic output and enabling the pricing of emissions-related risks in financial evaluations.
@@
Flow and Interaction Between Variables: The interaction between these variables provides a comprehensive framework for assessing the financial, operational, and environmental trade-offs involved in a low-carbon transition. For example, increasing the transition investment ratio ($\alpha$) may initially raise firm’s production costs ($c$), but this can lead to greater efficiency ($k$) and reduced carbon intensity ($I$) for the firm over time. Firms in markets with high carbon prices ($B$) are more likely to invest in low-carbon technologies, seeing long-term benefits from both regulatory compliance and enhanced profitability through premium pricing and consumer demand for sustainable products. By considering these variables, companies can strategically navigate the low-carbon transition, balancing short-term financial considerations with long-term environmental and competitive advantages. This paper will build the analytical framework based on these parameters and variables in the following parts.
In summary, the variables and associated parameters of the analytical framework are described in Table 1.
@@
$p-c$ Selling Profit per unit

### v2

@@ Abstract
This paper investigates strategic investments to mitigate transition risks in sectors mostly impacted by the shift to a low-carbon economy. It emphasizes the importance of tailored sector-specific strategies and the role of government interventions in shaping corporate behavior. Using a stylized, multi-period model, this paper evaluates the economic and operational trade-offs companies face under four decarbonization scenarios. The analysis provides practical insights for both policymakers and business leaders, demonstrating how regulatory frameworks and strategic investments can be aligned to manage transition risks while optimizing long-term sustainability.
@@ 1 Introduction
Mitigating climate change requires an earnest effort of decarbonization of the global economy. Achieving decarbonization demands transformative changes across sectors of industries in the economy. While sectors such as a power generation have made strides toward a renewable energy adoption, industries such as a steel and aviation lag behind in this effort due to underdeveloped low-carbon alternatives (Kober et al. (2020), Dietz et al. (2020)). These challenges highlight the importance of understanding climate risks, which fall into two categories: physical and transition risks. Physical risks encompass direct impacts of climate change, such as extreme weather events, while transition risks arise from shifting to a low-carbon economy driven by regulations, technologies, and societal demands (Carney (2015)). This paper focuses its analysis on transition risks as they offer a feasible scope for a strategic intervention and remain relatively underexplored in the field of quantitative modeling in sustainable finance (Cisagara (2024)).
@@
Transition risks result from structural changes in industries due to a shift toward a low-carbon economy. High-emission sectors, such as steel, aviation, and energy-intensive manufacturing, face challenges in adapting to stricter regulations and cleaner technologies. Policy uncertainty adds complexity, creating financial risks for firms and investors (Holden et al. (2018)). These risks include stranded assets, increased credit risk, and volatility in asset returns (Semieniuk et al. (2021)).
@@
While low-carbon policies, such as carbon taxes and emissions trading schemes, are essential for meeting climate objectives, their implementation introduces non-trivial transitional challenges to companies (Bank of England (2018)). Companies are evaluated not only on their financial metrics but also on their resilience to climate risks (Schulte and Hallstedt (2018)). Addressing these risks effectively requires targeted strategies and robust quantitative models.
@@
The shift toward a low-carbon economy offers opportunities to firms, such as cost savings, market leadership, and innovation, alongside major challenges. Companies like Tesla and BP have successfully aligned with low-carbon strategies, reaping benefits such as new revenue streams and enhanced competitiveness. However, this effort of transition is capital-intensive, with high upfront costs and regulatory uncertainty posing risks to companies (NGFS (2019), Schulte and Hallstedt (2018)). Additionally, the rapid pace of technological advancements can render current investments obsolete, further complicating any long-term planning that companies are undertaking.
@@
Policymakers and businesses must balance a climate action with financial stability. In the next section, we introduce a highly stylized model to guide low-carbon transition strategies, helping industries navigate uncertainties and optimize their transformational efforts.
2 Preliminary Steps toward a Quantitative Model
@@ 2.1 Motivations
Different industries face different and unique challenges and investment needs in the low-carbon transition. Key sectors in the economy include: (i) Energy, where transitioning to renewable energy requires upgrading infrastructure and advancing storage technologies (Boyd (2013)); (ii) Transportation, where decarbonization efforts focus on electric vehicles, hydrogen fuel, and charging networks (Tyfield (2013)); (iii) Manufacturing, where industries such as steel and chemical processing are exploring low-carbon processes such as hydrogen-based production and carbon capture technologies; and (iv) Agriculture, where investments in regenerative practices and ecosystem preservation are critical for reducing emissions (van Veelen (2021)).
Transition costs are substantial to firms. Estimates suggest that a global spending on net-zero initiatives could reach \$275 trillion by 2050, with varying investment levels across sectors (McKinsey (2021)). Therefore, tailored, sector-specific strategies are called for to balance costs and achieve decarbonization goals more effectively.
Investing in the low-carbon transition confers great benefits across various industries by driving long-term sustainability goal, reducing regulatory risks, and enhancing competitive positioning. Reducing carbon emissions can lead to cost savings through improved energy efficiency and reduced fuel expenses for energy-intensive industries such as manufacturing and transportation. In the financial sector, supporting low-carbon initiatives can mitigate portfolio risks linked to climate change, while at the same time also opening new investment opportunities in green technologies and sustainable businesses. Industries such as real estate and construction benefit by developing energy-efficient buildings that attract environmentally conscious tenants, which boost property values. Additionally, adopting low-carbon strategies can enhance the brand reputation for corporate responsibility and environmental stewardship, meet evolving consumer preferences for sustainability, and ensure compliance with future climate regulations, positioning companies for growth in a low-carbon economy. However, the upfront costs of adopting cleaner technologies can strain businesses, particularly if the returns on investment are not materialized immediately. For example, investing in renewable energy infrastructure may tie up capital that could otherwise be used for business expansion or other needs.
@@
Given these considerations, this paper proposes a highly stylized model to determine the optimal low-carbon transition strategy across different sectors of the indutries in the economy. The model aims at capturing the unique underlying dynamics and investment needs of each sector, helping industries navigate the complexities of transitioning toward a sustainable, low-carbon future.
2.2 Components of the Model
@@
Variables are introduced in this model to replicate the dynamics of low-carbon transition investments and their effects on various industry operations, offering insights into investment strategies. Below, we provide a detailed description of each of the variables and parameters of the model.
@@
Selling price per unit ($p$). This variable represents the price at which the product of interest is sold in the market. Taking the selling price into consideration is to describe and quantify the revenue generated by the firm. In the context of low-carbon transition, businesses may find opportunities to increase prices for eco-friendly products due to an increased consumer demand for sustainability. For instance, Tesla’s Model 3 and other EVs are often priced higher than traditional internal combustion engine vehicles (ICEs), yet some consumers are still willing to pay the premiums for the environmental advantages.
@@
Production Cost per unit ($c$). This variable includes all expenses incurred by a company in manufacturing each unit, such as raw materials, labor, and energy. Analyzing production costs is essential to evaluating profitability of the company. A study by Siddique and Sultana (2018) found that switching to sustainable sourcing practices initially raised production costs for the company due to the adoption of cleaner and more expensive technologies, but long-term benefits such as energy savings and brand value will eventually outweigh the initial investment.
@@
Selling Profit per unit (pcp-c). This variable reflects the profit earned from each unit sold by a company. It is determined by subtracting the production costs from the selling price. It captures the direct profitability of the product after accounting for costs. Profit is the primary goal for a company when making decisions, as it ensures financial viability and supports long-term growth, influencing how much of the company’s revenue can be allocated toward investments, including those for sustainable transitions.
@@
Total Asset ($A$). This variable represents all financial resources available to a company, encompassing both current and long-term assets. This metric is essential for evaluating the company’s capacity to invest in production and growth opportunities, providing insight into the scale and potential of the business. Research indicates that firms with larger asset bases are more inclined to invest more substantially in sustainability initiatives, as they can support longer investment horizons and absorb higher upfront costs associated with such projects (Eccles et al. (2014)). For example, Amazon has invested \$2 billion in renewable energy through its Climate Pledge Fund, leveraging its extensive financial resources to lead in corporate sustainability (Amazon (2020)). Consequently, total assets not only reflect a company’s financial strength but also its commitment to sustainable practices and long-term value creation.
@@
Transition Investment Ratio ($\alpha$). This variable is a critical metric in our model. It measures the proportion of total assets allocated to low-carbon initiatives. This ratio is essential to a firm as it reflects its commitment to sustainability and its responsiveness to market dynamics and regulatory pressures. By strategically adjusting this ratio, businesses can enhance their reputation and mitigate risks associated with climate-related regulations and consumer preferences. Our primary goal of working with this model is to identify the optimal transition investment ratio that balances financial performance with sustainability objectives, ensuring that companies thrive economically while they are still able to contribute positively to the environment.
@@
In practice, the Transition Investment Ratio offers insights into how effectively a company aligns its financial resources with its sustainability goals. A higher value of this ratio signifies a greater investment in low-carbon initiatives, leading to long-term cost savings and competitive advantages. For instance, IKEA’s allocation of over 3 billion pounds toward a renewable energy infrastructure exemplifies the benefits of a strategic adjustment in this ratio (IKEA (2020)). By investing in solar and wind energy projects, IKEA not only reduces its operational reliance on fossil fuels but also achieves improved operational efficiency and lower carbon emissions. Improved operational efficiency can manifest in reduced energy costs, as renewable energy sources often lead to lower utility bills. IKEA’s renewable energy infrastructure allows it to produce a substantial portion of its energy needs in a sustainable manner, ultimately lowering the company’s operational costs and enhancing its energy independence. Additionally, the transition to renewable energy directly impacts carbon emissions, as utilizing solar and wind energy decreases IKEA’s greenhouse gas footprint. This shift aligns with regulatory expectations and resonates with consumers who favor environmentally responsible brands. Thus, the Transition Investment Ratio plays a pivotal role in our model as it encapsulates the strategic allocation of resources toward sustainability and driving both operational efficiency and lower carbon emissions.
@@
As previously mentioned, undergoing a low-carbon transition and investing in this transition can decrease production costs and increase selling prices for the company due to improved operational efficiency, stronger brand value, and an alignment with consumer preferences. In the model, we assume both the production costs and the selling prices to be constant for simplicity. This assumption allows us to isolate the decision-making process regarding the resource allocation and emphasize the role of the Transition Investment Ratio in balancing financial performance and sustainability objectives without introducing additional complexities.
@@
Enterprise Low-carbon Production Efficiency Coefficient ($k$). This coefficient measures the improvement in the firm’s production efficiency resulting from low-carbon initiatives, with a value greater than 1 indicating a gain in efficiency due to cleaner technologies. In our model, we assume that a higher value of $k$ indicates greater production efficiency of the firm after investing in the transition to sustainable practices. For instance, Siemens reported a remarkable 20% increase in production efficiency after transitioning from traditional fossil fuel-based energy to renewable energy sources and incorporating energy-efficient manufacturing technologies (AG (2019)). This transition not only reduced their carbon footprint but also streamlined their operations, demonstrating that investments in sustainability can yield great returns in efficiency.
@@
Original Productivity Coefficient ($\beta$). This coefficient is set to range between 0 and 1. A lower value indicates that, given a fixed amount of raw materials, the firm’s production efficiency is relatively low, resulting in fewer products being produced per unit of input, which reflects suboptimal resource utilization. Conversely, a value closer to 1 suggests higher efficiency in material utilization, allowing the enterprise to produce more output with the same raw material input, indicating both optimized resource allocation and production process efficiency. Changes in this coefficient can thus reflect the current production efficiency of the enterprise and offer insights into potential improvements under low-carbon transition strategies. There is some evidence to suggest that many firms that start with below-average efficiency have successfully turned their operations around after investing in cleaner technologies. Notably, research in Weng et al. (2015) indicates that companies in the regions with lower initial productivity often experience the most substantial gains when transitioning to more efficient, low-carbon systems. By enhancing productivity through sustainable practices, these firms not only are able to improve their operational performance but they also are able to position themselves for greater profitability in the long run. Thus, the firms with lower $\beta$ values should prioritize upgrades to enhance productivity and better align with sustainability goals.
@@
Carbon Price ($B$). This variable captures the cost of emissions imposed by regulatory frameworks and is essential to a company’s profitability. This price is sensitive to regulatory or policy changes, with high carbon prices creating a financial pressure on firms to reduce emissions. When the market carbon price is elevated, it can diminish a company’s profit margins, compelling businesses to reassess their operational strategies. In response to these challenges, companies are incentivized to invest in cleaner technologies and adopt sustainable practices to mitigate their carbon liabilities. As firms strive to maintain profitability while adhering to environmental regulations, understanding and managing the implications of carbon pricing becomes crucial for their long-term success.
@@
Carbon Intensity ($I$). This variable refers to the amount of carbon dioxide (CO2) emissions produced per unit of output. It is typically measured in terms of emissions per product manufactured or per unit of energy consumed. It serves as a critical metric for companies transitioning to low-carbon production, as reducing carbon intensity is essential for minimizing environmental impact while maintaining operational efficiency.
@@
In the literature, carbon intensity is usually defined as a measure of the environmental performance of a company’s production processes. For instance, the Intergovernmental Panel on Climate Change (IPCC) emphasizes that carbon intensity is a key indicator of how effectively an organization utilizes resources and minimizes emissions in its operations (Change et al. (2014)). Similarly, the World Resources Institute (WRI) defines carbon intensity as a means to assess the relationship between output and emissions, allowing for comparisons across industries and over time (Institute (2015)). In addition, carbon intensity is defined more specifically as the ratio of expected annual emissions to expected revenue in Fang et al. (2019). This formulation highlights the firm’s contribution to climate change per unit of revenue, thus linking environmental impact directly to economic output and enabling the pricing of emissions-related risks in financial evaluations.
@@
Flow and Interaction between Variables. The interaction between these variables allows us to assess the financial, operational, and environmental trade-offs involved in a low-carbon transition for companies. For example, increasing the transition investment ratio ($\alpha$) may initially raise firm’s production costs ($c$), but this can lead to greater efficiency ($k$) and reduced carbon intensity ($I$) for the firm over time. Firms in markets with high carbon prices ($B$) are more likely to invest in low-carbon technologies, seeing long-term benefits from both regulatory compliance and enhanced profitability through premium pricing and consumer demand for sustainable products. By considering these variables, companies can strategically navigate the low-carbon transition, balancing short-term financial considerations with long-term environmental and competitive advantages. This paper constructs a highly stylized, quantitative model based on these variables.
The variables of our proposed model are summarized in Table 1.
@@
pcp-c Selling Profit per unit